# Are Users More Willing to Use Formally Verified Password Managers?

### v1

@@ Abstract.
Formal verification has recently been increasingly used to prove the correctness and security of many applications. It is attractive because it can prove the absence of errors with the same certainty as mathematicians proving theorems. However, while most security experts recognize the value of formal verification, the views of non-technical users on this topic are unknown. To address this issue, we designed and implemented two experiments to understand how formal verification impacts users. Our approach started with a formative study involving 15 participants, followed by the main quantitative study with 200 individuals. We focus on the application domain of password managers since it has been documented that the lack of trust in password managers might lead to lower adoption. Moreover, recent efforts have focused on formally verifying (parts of) password managers. We conclude that formal verification is seen as desirable by users and identify three actional recommendations to improve formal verification communication efforts.
@@
Password Manager, Usable Security, Formal Methods, Formal Verification, User Study, HCI
journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Usability in security and privacyccs: Security and privacy Formal methods and theory of security
1. Introduction
@@
Formal verification has developed substantially in recent years and has been applied to many domains. Recent work in this area includes the formal verification of security properties in trusted execution environments (jangid2021towards, ), critical parts of the Linux kernel (nelson2020specification, ), the open-source TLS implementation used in numerous Amazon services (chudnov2018continuous, ), and even compilers (leroy2006formal, ; kumar2014cakeml, ). Formal verification works by rigorously proving the correctness of code against a formal specification and thus ensuring that software components behave as intended, even under adversarial conditions. An extensive survey conducted in 2020 (2020_expert_surveyFM, ) revealed that experts in formal methods widely believe these approaches offer enhanced code quality, strengthened cybersecurity, streamlined certification processes, and more manageable maintenance efforts. Although security experts generally recognize the value of formal verification, the views (or even awareness) of end-users are unknown (carreira2021exploring, ; carreira2022studying, ).
@@
Previous research has explored user expectations and perceptions of security in areas such as differential privacy (cummings2021need, ) and cryptocurrency systems (mai2020user, ). Similarly, prior work on formal methods has also successfully evidenced trends, limitations, and future paths for formal methods (2020_expert_surveyFM, ). However, to the best of our knowledge, the impact of formal verification on end-user willingness to use formally verified products is still unexplored. Understanding users’ views on formal verification can help foster the adoption of formally verified code (with all the advantages verified code can provide) and aid developers in communicating with end-users and industry partners.
@@
Our work addresses this gap in the literature by studying formal verification’s impact on users of a concrete application domain: Password Managers. Password Managers are applications that help users generate and manage their passwords. Most Password Managers offer password security features like password generation, multi-factor authentication, and secure storage with a primary password. [^1] Some related work (pearman2019people, ; ray2020older, ; chiasson2006usability, ) uses “master password” instead of “primary password”. However, recent work uses “primary password” (e.g., Firefox’s Password Manager (mozilla, )).
@@
We chose this domain because text passwords are one of the most used security mechanisms, and Password Managers are an essential security tool to manage them (persistenceOfPasswords, ). However, previous work suggests that users have trust issues with Password Managers (ion2015no, ; ray2020older, ) and are reluctant to use them due to a lack of understanding of their security properties (pearman2019people, ). In parallel, recent work on formal verification has been applied to Password Managers (grilo2021towards, ; grilo22pwdgen, ). We designed and deployed two studies to understand users’ views on formal verification. The first is a formative interview study, asking users to use a Password Manager with verified components. The main study is a more extensive quantitative study that builds upon the formative and aims to explore some of the findings and themes identified in more detail.
@@
Addressing a critical gap in the literature: Our research provides an empirical evaluation of the influence of formal verification on end-users within a specific application domain —- Password Managers. This work makes a contribution by addressing a gap in the formal methods literature – the study of end-users of formal verification.
Impact of formal verification on user adoption: Our findings suggest that formal verification has a positive effect on users’ willingness to adopt Password Managers. In both our formative user study and main study, participants showed an increased willingness to use Password Managers with formally verified components and identified formal verification as a desirable feature.
Identification of features for formal verification: We systematically identify a hierarchy of Password Manager features that users consider most critical for formal verification. We identify several critical features that are particularly relevant for participants. We argue that developers and practitioners should focus their formal verification efforts on features that users deem most critical for their trust and usage intentions, such as vault security and password generation.
@@
2. Scope and Research Questions
Our formative study motivates three questions. This study aims to test users’ perceptions of formal verification. To this end, we choose the domain of Password Managers and specifically address the following Research Questions:
RQ1.:
How does formal verification impact users’ willingness to use Password Managers?
RQ2.:
What features would users like to see formally verified in a Password Manager?
RQ3.:
Do users value the guarantees formal verification can provide in Password Managers?
@@ 2. Scope and Research Questions
Figure 1. Overview of the Methodology
Our overarching goal is to understand how formal verification in Password Managers influences the decisions and trust of non-expert users.
The formative study, characterized by its qualitative and exploratory nature, focuses solely on RQ1. It aims to capture users’ initial reactions and understandings of formal verification in Password Managers through in-depth interviews and interactions with a Password Manager prototype. This prototype, an extended version of Bitwarden (bitwarden, ), incorporates a formal verification icon and explanations to educate users about formal verification’s role in enhancing Password Manager security. It gathers insights on general themes around Password Managers and formal verification, setting the groundwork for the subsequent larger-scale survey. The emphasis here is on identifying and exploring users’ perceptions, which directly contributes to addressing RQ1 by revealing how users view and understand formal verification in the context of Password Managers.
The main study builds on the findings from the formative study, expanding the scope to address all research questions mentioned in the introduction. With 200 participants, this larger-scale survey aims to quantitatively assess the broader implications of formal verification on users’ willingness to use Password Managers. This study aims to validate the themes identified in the formative study at a larger scale and further explore how formal verification impacts users’ choices and trust in Password Managers. We use a combination of statistical analysis and scenario-based questions to understand better the role of formal verification across different aspects of Password Manager usage and answer our research questions.
Crucially, we do not wish to ”explain” formal verification to our participants; we merely want to communicate its consequences and impact on the Password Manager. An overview of the methodology adopted can be seen in Figure 1.
@@
In both studies, we did not collect any personal data. All participants were over 18 years old and were shown a consent form, the terms of which they had to accept before starting the survey. In the first author’s institution (the lead institution for this project), studies with these properties do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee). The local Ethics Committee recommends reviewing the information and checklists shown in the European Commission’s guide on ethical self-assessment. [^2] European Commission’s guide on ethical self-assessment can be accessed here https://ec.europa.eu/info/funding-tenders/opportunities/docs/2021-2027/common/guidance/how-to-complete-your-ethics-self-assessment_en.pdf. All studies described in this paper followed these guidelines. Co-authors from other institutions did not have access to individual survey data and only worked with aggregate statistics, as reported in this paper.
3. Background and Motivation
@@
In general, it is hard to build secure computer-based systems. Formal methods offer the promise of software that does not have exploitable bugs (fisher2017hacms, ) and has been used in a wide range of domains to prove the correctness and security of applications (jangid2021towards, ; nelson2020specification, ; grilo22pwdgen, ; chudnov2018continuous, ; ferreira2017certified, ; grilo2022verified, ). The idea of formal verification is mathematically modeling a system and then using formal methods to prove that the model satisfies specific properties or specifications. For example, Chudnov et al. (chudnov2018continuous, ) describe the development and operation of a continuously checked proof ensuring key properties of the TLS implementation used by many AWS services. Nelson et al. (nelson2020specification, ) describe their experience applying formal methods to a critical component in the Linux kernel, the just-in-time compilers. Their results show that building a verified component within a sizeable unverified system is possible with careful specification and proof strategy design. Another line of work investigates the application of formal methods to the specification and evaluation of password composition policies. Johnson et al. (johnson2020skeptic, ) introduce Skeptic, a toolchain that applies formal verification and power-law modeling to automatically select and justify password composition policies. Formally verified compilers are also available. An example is the CompCert (leroy2009formal, ) compiler, which compiles code from a large subset of the C programming language to PowerPC assembly code and guarantees that safety properties proved on the source code hold for the executable compiled code as well (leroy2009formal, ).
Password managers.
Passwords have become increasingly complex as most websites adopt strict password security policies.Previous work suggests users struggle with generating unique passwords (inglesant2010true, ), and password reuse is a problem—Pearman et al. (pearman2017let, ) show that 40% of users reuse 81–90% of their passwords. The basic features that all Password Managers provide to users are: (a) A secure encrypted vault to save passwords; (b) A password generator to generate unpredictable passwords that comply with complex password security policies. Moreover, most modern Password Managers also provide other features, including: (c) “Autofill” to automatically fill passwords in a website’s login screen; (d) Cloud synchronization to keep passwords safe and synchronized across devices. Despite being recommended by experts  (enisapassword, ), Password Manager usage is not widespread. Users complain about not knowing that Password Managers exist and not understanding how they work (pearman2019people, ; ion2015no, ; ray2020older, ; chiasson2006usability, ). Users also do not trust Password Managers (pearman2019people, ). As an intermediary tool, Password Managers must be reliable, consistent, and predictable. Previous work (chiasson2006usability, ; pearman2019people, ) recommends education as a tool for enhancing adoption and trust in Password Managers.
@@
An effective security mechanism is one that is used correctly. Usable security can be traced back to Saltzer and Schroeder’s 1974 paper (saltzer1975protection, ) that introduces the term “psychological acceptability” for access-control systems. In the last decade, there has been a lot of work on usable security across many domains (mai2020user, ; pearman2019people, ; ion2015no, ; carreira2021towards, ). Prior work explored various dimensions of security warnings, from SSL warnings in web browsers to end-to-end encryption guarantees. A notable contribution is the series of studies by Felt et al. (felt2015improving, ) examining the impact of SSL warnings on user behavior within Google Chrome. Their work focused on experimenting with SSL warnings and assessing user adherence and comprehension. Similarly, Akhawe and Felt’s (akhawe2013alice, ) study provided evidence on the effectiveness of browser security warnings, revealing insights into how users perceive and react to these alerts in real-world settings. Other relevant, usable security work includes previous efforts to communicate about security using labels similar to “nutrition labels (kelley2009nutrition, ).” In this realm, previous research designed privacy labels for several domains, such as app stores (li2022understanding, ) and IoT devices (emami2020ask, ). Related work has also evaluated the effectiveness of commonly deployed password strength meters and found significant inconsistencies between meter classifications and actual password resistance to cracking: many passwords labeled as “strong” were, in fact, easily guessed, and some “weak” passwords resisted cracking (pereira2020evaluating, ).
@@
A 2020 survey (2020_expert_surveyFM, ) of formal methods experts studied their views of formal methods. However, to the best of our knowledge, no previous work has focused on (non-expert) users of formally verified Password Managers, nor, more generally, on the impact that formal verification may have on end users. Previous work on formal verification of Password Managers is scarce, but efforts have been made to formally verify the random password generation algorithm of a Password Manager (grilo2021towards, ; grilo22pwdgen, ). However, these studies concentrated on the technical aspects of formal verification without examining its usability implications.
4. Formative Study
Our goal with this formative study was to understand users’ overall views on formal verification. To achieve this, we implemented two significant extensions to a Password Manager: a formal verification icon and an explanation.
4.1. Method
We created a proof-of-concept prototype by extending an established Password Manager’s – Bitwarden’s (bitwarden, ) browser extension – interface and testing it for user acceptance.
@@ 4.1. Method
Bitwarden uses icons to communicate with users. Similarly, we use an icon to communicate formal verification as icons are a good alternative to text to communicate with users (Wiedenbeck1999, ). The formal verification icon we chose is shown in Figure 2 and was placed on all features that could be formally verified: (a) Password vault, by the password field; (b) Primary password input box; (c) Password generator; (d) Clipboard, by its settings.
We chose green due to its association with safety in Anglo-American contexts (dix2004human, ) and derived the green used in the UI from Bitwarden’s blue (#175DDC ) and grey (#7C7C7C ), ending up with the shade (#0BDB0B ).
We designed the formal verification icon with Font Awesome 4333Font Awesome is fully open source and GPL font friendly. https://fontawesome.com, the same font used to design the existing Bitwarden’s iconography. We chose green due to its association with safety in Anglo-American contexts (dix2004human, ) and derived it from Bitwarden’s blue (#175DDC) and grey (#7C7C7C), ending up with the shade (#0BDB0B) using Adobe Color [^4] A palette tool referenced by Tidwell et al. (tidwell2010designing, ). https://color.adobe.com/create. The design process followed three phases: (1) Initial brainstorming with security researchers that resulted in 15 icon designs, later narrowed to 7. (2) Contextual tests of the icons within the PMs interface led to feedback on aesthetics, size, and positioning. (3) An external survey with 20 users with: (a) an initial preference test without explanation, (b) a preference test informed by the icon’s purpose (harley_2016, ), (c) final ratings in the icon’s intended context. A detailed description of the icon design process can be found in Appendix A.
Figure 2. Formal Verification icon
@@
To explain the role of formal verification in the Password Manager, we initially identified all features amenable to formal verification, as well as the potential locations within the Password Manager interface for displaying the verification icon.
We developed the explanations in three phases: (1) We began with one-on-one discussions with formal verification researchers, gathering explanations, and finally, removing jargon. (2) The second iteration consisted of meeting with the team, gathering feedback , and applying it. (3) Finally, after another round of feedback, we reached a consensus and finalized the explanations. We compared the base Password Manager (i.e., without any extensions or formal verification) and the extended Password Manager (i.e., with interface updates). We performed two pilot interviews to refine the protocol and interview script.
@@
The user tests were divided into four parts. First, we provided users with a brief introduction to the study, and after we asked users to fill out the “Pre-Task Questionnaire” about their experience with Password Managers and demographics. All participants were asked to perform everyday tasks in a Password Manager (e.g., save a password). And finally, we ask them to fill out the “Final Questionnaire” . Each questionnaire was used as a base for semi-structured interviews. We asked follow-up questions about the participants’ answers. Data was collected in two ways: through questionnaires and observation. The extended Password Manager did not have formally verified features. As such, we included the following disclaimer: “This is a product in development, and some formally verified features are not fully implemented.”.
@@
Our sample comprised 15 participants – 10 for the extended interface and 5 for the control condition. Participants were recruited through the authors’ network and received no compensation. Most of the participants (60%) had higher education. Of the 15 participants, only 2 had a technical background related to IT. The most frequent age group was 25-34 (40%). Overall, 60% of users were younger than 34, and 40% were at least 35 years old. Gender in our sample was evenly divided. Each interview took around 50 minutes, and 90% of participants were unfamiliar with formal verification.
4.2. Results
To understand if participants knew the icon’s (see Figure 2) meaning, we asked them to explain it. Of the 10 participants, 5 mentioned formal verification, and the other five mentioned concepts related to the security of the Password Manager (e.g., “the icon means that the passwords were safe” (P1)).

### v2

@@ Abstract
Formal verification has recently been increasingly used to prove the correctness and security of many applications. It is attractive because it can prove the absence of errors with the same certainty as mathematicians proving theorems. However, while most security experts recognize the value of formal verification, the views of non-technical users on this topic are unknown. We designed and implemented two experiments to address this issue to understand how formal verification impacts users. Our approach started with a formative study involving 15 participants, followed by the main quantitative study with 200 individuals. We focus on the application domain of Password Managers (PMs) since it has been documented that the lack of trust in PMs might lead to lower adoption. Moreover, recent efforts have focused on formally verifying (parts of) PMs. We conclude that formal verification is seen as desirable by users and identify three actionable recommendations to improve formal verification communication efforts.
Keywords:
@@
Password Managers, Usable Security, Formal Methods, Formal Verification, User Study, HCI, User Trust, Security Adoption, Verified Software, Software Assurance, Human Factors in Security, Cryptography Usability, Secure Software Adoption
1 Introduction
@@
Formal verification has developed substantially in recent years and has been applied to many domains. Recent work in this area includes the formal verification of security properties in trusted execution environments [27], critical parts of the Linux kernel [40], the open-source TLS implementation used in numerous Amazon services [10], and even compilers [34, 31]. Formal verification works by rigorously proving the correctness of code against a formal specification and thus ensuring that software components behave as intended, even under adversarial conditions. An extensive survey conducted in 2020 [18] revealed that experts in formal methods widely believe these approaches offer enhanced code quality, strengthened cybersecurity, and more manageable maintenance efforts. Although security experts generally recognize the value of formal verification, the views (or even awareness) of end-users are unknown [8, 6].
@@
Previous research has explored user expectations and perceptions of security in areas such as differential privacy [11] and cryptocurrency systems [37]. Similarly, prior work on formal methods has also successfully evidenced trends, limitations, and future paths for formal methods [18]. However, to the best of our knowledge, the impact of formal verification on end-user willingness to use formally verified products is still unexplored. Understanding users’ views on formal verification can help foster the adoption of formally verified code (with all the advantages verified code can provide) and aid developers in communicating with end-users and industry partners.
@@
Our work addresses this gap in the literature by studying formal verification’s impact on users of a concrete application domain: Password Managers (PMs). PMs are applications that help users generate and manage their passwords. Most PMs offer password security features like password generation, multi-factor authentication, and secure storage with a primary password.
@@
We chose this domain because text passwords are one of the most used security mechanisms, and PMs are an essential security tool to manage them. However, previous work suggests that users have trust issues with PMs [26, 45] and are reluctant to use them due to a lack of understanding of their security properties [41]. In parallel, recent work on formal verification has been applied to PMs [20]. We designed and deployed two studies to understand users’ views on formal verification. The first is a formative interview study. The main study is a more extensive quantitative study that builds upon the formative and aims to explore some of its findings in more detail.
@@
Addressing a critical gap in the literature: Our research provides an empirical evaluation of the influence of formal verification on end-users within a specific application domain — PMs. This work makes a contribution by addressing a gap in the formal methods literature — the study of end-users of formally verified products.
Impact of formal verification on user adoption: Our findings suggest that formal verification has a positive effect on users’ willingness to adopt PMs. In both our formative user study and main study, participants showed an increased willingness to use PMs with formally verified components and identified formal verification as a desirable feature.
Identification of features for formal verification: We systematically identify a hierarchy of PM features that users consider most critical for formal verification. We identify several critical features that are particularly relevant for participants. We argue that developers and practitioners should focus their formal verification efforts on features that users deem most critical for their trust and usage intentions, such as vault security and password generation.
@@
2 Scope and Research Questions
Our study aims to test users’ perceptions of formal verification. To this end, we choose the domain of PMs and address the following Research Questions:
RQ1.
How does formal verification impact users’ willingness to use PMs?
RQ2.
What features would users like to see formally verified in a PM?
RQ3.
Do users value the guarantßees formal verification can provide in PMs?
@@ 2 Scope and Research Questions
Our overarching goal is to understand how formal verification in PMs influences the decisions and trust of non-expert users.
The formative study, characterized by its qualitative and exploratory nature, focuses solely on RQ1. It aims to capture users’ initial reactions and understandings of formal verification in PMs through in-depth interviews and interactions with a PM prototype. This prototype, an extended version of Bitwarden [5] incorporates a formal verification icon and explanations to educate users about formal verification’s role in enhancing PM security. It gathers insights on general themes around PMs and formal verification, setting the groundwork for the subsequent larger-scale survey. The emphasis here is on identifying and exploring users’ perceptions, which directly contributes to addressing RQ1 by revealing how users view and understand formal verification in the context of PMs.
The main study builds on the findings from the formative study, expanding the scope to address all RQs mentioned above. With 200 participants, this larger-scale survey aims to quantitatively assess the broader implications of formal verification on users’ willingness to use PMs. This study aims to validate the themes identified in the formative study at a larger scale and further explore how formal verification impacts users’ choices and trust in PMs. We use a combination of statistical analysis and scenario-based questions to better understand the role of formal verification across different aspects of PM usage and answer our RQs.
Crucially, we do not wish to “explain” formal verification to our participants; we merely want to communicate its consequences and impact on the PM. An overview of the methodology adopted can be seen in Figure˜1.
@@
In both studies, we did not collect any personal data. All participants were over 18 years old and were shown a consent form, which they had to accept before starting the survey. In the first author’s institution (the lead institution for this project), studies with these properties do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee). The local Ethics Committee recommends reviewing the information and checklists shown in the European Commission’s guide on ethical self-assessment. [^1] European Commission’s guide on ethical self-assessment can be accessed here https://ec.europa.eu/info/funding-tenders/opportunities/docs/2021-2027/common/guidance/how-to-complete-your-ethics-self-assessment_en.pdf. All studies described in this paper followed these guidelines.
Figure 1: Overview of the Methodology
3 Background and Motivation
@@
In general, it is hard to build secure computer-based systems. Formal methods offer the promise of software that does not have exploitable bugs and has been used in a wide range of domains to prove the correctness and security of applications [27, 40, 20, 10, 16]. The idea of formal verification is mathematically modeling a system and then using formal methods to prove that the model satisfies specific properties or specifications. For example, Chudnov et al. [10] describe the development and operation of a continuously checked proof ensuring key properties of the TLS implementation used by many AWS services. Nelson et al. [40] describe their experience applying formal methods to a critical component in the Linux kernel, the just-in-time compilers. Their results show that building a verified component within a sizeable unverified system is possible with careful specification and proof strategy design. Another line of work investigates the application of formal methods to the specification and evaluation of password composition policies. Johnson et al. [28] introduce Skeptic, a toolchain that applies formal verification and power-law modeling to automatically select and justify password composition policies. Formally verified compilers are also available. An example is the CompCert [35] compiler, which compiles code from a large subset of the C programming language to PowerPC assembly code and guarantees that safety properties proved on the source code hold for the executable compiled code as well [35].
PMs.
Passwords have become increasingly complex as most websites adopt strict password security policies. Previous work suggests users struggle with generating unique passwords [25], and password reuse is a problem — Pearman et al. [42] show that 40% of users reuse 81–90% of their passwords. The basic features that all PMs provide to users are: (a) A secure encrypted vault to save passwords; (b) A password generator to generate unpredictable passwords that comply with complex password security policies. Moreover, most modern PMs also provide other features, including: (c) “Autofill” to automatically fill passwords in a website’s login screen; (d) Cloud synchronization to keep passwords safe and synchronized across devices. Despite being recommended by experts  [14], PM usage is not widespread. Users complain about not knowing that PMs exist and not understanding how they work [41, 26, 45, 9]. Users also do not trust PMs [41]. As an intermediary tool, PMs must be reliable, consistent, and predictable. Previous work [9, 41] recommends education as a tool for enhancing adoption and trust in PMs.
@@
An effective security mechanism is one that is used correctly. Usable security can be traced back to Saltzer and Schroeder’s 1974 paper [48] that introduces the term “psychological acceptability” for access-control systems.
In the last decade, there has been a lot of work on usable security across many domains [37, 41, 26, 7]. Prior work explored various dimensions of security warnings, from SSL warnings in web browsers to end-to-end encryption guarantees. A notable contribution is the series of studies by Felt et al. [15] examining the impact of SSL warnings on user behavior within Google Chrome. Their work focused on experimenting with SSL warnings and assessing user adherence and comprehension. Similarly, Akhawe and Felt’s [2] study provided evidence on the effectiveness of browser security warnings, revealing insights into how users perceive and react to these alerts in real-world settings. Other relevant, usable security work includes previous efforts to communicate about security using labels similar to “nutrition labels [30].” In this realm, previous research designed privacy labels for several domains, such as app stores [36] and IoT devices [13]. Related work has also evaluated the effectiveness of commonly deployed password strength meters and found significant inconsistencies between meter classifications and actual password resistance to cracking: many passwords labeled as “strong” were, in fact, easily guessed, and some “weak” passwords resisted cracking [43].
@@
A 2020 survey [18] of formal methods experts studied their views of formal methods. However, to the best of our knowledge, no previous work has focused on (non-expert) users of formally verified PMs, nor, more generally, on the impact that formal verification may have on end users. Previous work on formal verification of PMs is scarce, but efforts have been made to formally verify the random password generation algorithm of a PM [20] or look at usability [6]. However, these studies concentrated on the technical aspects of formal verification without examining its usability implications or had a smaller sample size.
4 Formative Study
Our goal with this formative study was to understand users’ overall views on formal verification. To achieve this, we implemented two significant extensions to a PM: a formal verification icon and an explanation.
4.1 Method
We created a proof-of-concept prototype by extending an established PM’s — Bitwarden’s browser extension — interface and testing it for user acceptance.
@@ 4.1 Method
We designed an icon to communicate formal verification ( 4 in  Figure˜1), and placed it on all features that could be formally verified: (a) Password vault; (b) Primary password input box; (c) Password generator; (d) Clipboard, by its settings. We derived the color green used in the UI from Bitwarden’s blue (#175DDC ) and grey (#7C7C7C ), ending up with the shade (#0BDB0B ). The design process followed three phases: (a) Initial brainstorming with security researchers that resulted in 15 icon designs, later narrowed to 7 ( 2 in Figure˜1). (b) Contextual tests of the icons within the PMs interface. (c) A survey with 20 users with two preference tests [22] and ratings in the icon’s intended context.
@@
To explain the role of formal verification in the PM, we began by identifying all verifiable features. We developed the explanations in three phases: (1) We began with one-on-one discussions with formal verification researchers, gathering explanations, and removing jargon ( 5 in Figure˜1). (2) We then met with the team, gathered feedback, and applied it. (3) Finally, after another round of feedback, we reached a consensus and finalized the explanations. We compared the base PM (i.e., without any extensions or formal verification) and the extended PM (i.e., with interface updates).
@@
The user tests were divided into the following parts. First, we provided users with a brief introduction to the study with a “Pre-Task Questionnaire” about their experience with PMs and demographics. We asked participants to perform everyday tasks in a PM (e.g., save a password). And finally, we asked them to fill out the “Final Questionnaire”. Each questionnaire was used as a base for semi-structured interviews. We collected data through questionnaires and observation.
@@
Our sample comprised 15 participants, 10 for the extended interface and 5 for standard interface. Participants were recruited through the authors’ network and received no compensation. Most of the participants (60%) had higher education. Of the 15 participants, only 2 had a technical background related to IT. The most frequent age group was 25-34 (40%). Overall, 60% of users were younger than 34. Gender in our sample was evenly divided. Each interview took around 50 minutes, and 90% of participants were unfamiliar with formal verification. We performed two pilot interviews to refine the protocol and interview script ( 6 in Figure˜1).
4.2 Results
To understand if participants knew the icon’s meaning (see  4 in Figure˜1), we asked them to explain it. Of the 10 participants, 5 mentioned formal verification, and the other five mentioned concepts related to the security of the PM (e.g., “the icon means that the passwords were safe” (P1)).